OpenStudy (anonymous):

Let P(x)=x^3-6x+k. If the sum of the two roots of P(x) is -2, then the third root is...

OpenStudy (anonymous):

r1 + r2 = -2, so you would say r2 = -r1 - 2 Then you have (x-r1)(x+r1+2)(x-r)

OpenStudy (anonymous):

where r is 3rd root.

OpenStudy (anonymous):

Should I be multiplying this out?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

no short cuts?

OpenStudy (anonymous):

Wolfram Alpha

OpenStudy (anonymous):

ye gods. Doing this during a test.

OpenStudy (anonymous):

You could so (x-a)(x-b)(x-c) and then do the b = -a-2 substitution after you are done.

OpenStudy (anonymous):

do^

OpenStudy (anonymous):

I'm halfway already. Canceling things out.

OpenStudy (anonymous):

x^3+(-r1-r+2)x^2+2(-r1+1)x+r1r(2+r1)

OpenStudy (anonymous):

You have three equations now. (-r1-r+2) = 0 2(-r1+1) = -6 r1r(2+r1) = k

OpenStudy (anonymous):

Solving this makes me very happy, but are you /absolutely/ sure there isn't a shorter way?

OpenStudy (anonymous):

There could be a shorter way, but I don't know it. I like to use the basics to do these things.

OpenStudy (anonymous):

Alrighty.

OpenStudy (anonymous):

I found the better solution: Vieta's Formula Let P(x)=x^3-6x+k. If the sum of the two roots of P(x) is -2, then the third root is... r1r2r3=k r1r2+r1r3+r2r3=-6 r1+r2+r3=0 r1+r2=-2 -2+r3=0 r3=2